OpenStudy (anonymous):

how do you multiply 3/1 by 1/2?

OpenStudy (lgbasallote):

\[\frac 31 \times \frac 12 \implies \frac{3 \times 1}{1 \times 2}\] does that help?

OpenStudy (anonymous):

I believe so :)

OpenStudy (lgbasallote):

wonderful

OpenStudy (anonymous):

thank you

OpenStudy (lgbasallote):

welcome